# Searches for long-lived dark photons at proton accelerator experiments

### v2

@@ Abstract
A systematic and unified study of the ability of lifetime frontier experiments to explore the parameter space of hypothetical long-lived particles is one of the main steps in defining their parameter space. Such an analysis has not been conducted for dark photons – hypothetical massive particles that have kinetic mixing with Standard Model photons. Existing studies have varied in their assumptions about dark photon phenomenology, often using outdated models that do not reflect recent advancements. In this paper, we present a unified calculation of the parameter space for GeV-scale dark photons probed by lifetime frontier experiments, delineating the regions excluded by past experiments and those accessible to future experiments. Our approach incorporates the latest advances in studying production mechanisms of dark photons with mass $m_{V}\gtrsim 0.1\text{ GeV}$, including proton bremsstrahlung and mixing with neutral mesons, and utilizes the full palette of hadronic decays. Additionally, we explore the impact of uncertainties in proton bremsstrahlung on the dark photon parameter space, and find that they may severely affect the reach of many past and future experiments, including the maximal probed mass. The results are provided in a publicly accessible format, specifically through the implementation of the updated phenomenological models in SensCalc – a unified tool for calculating event rates of new physics particles at lifetime frontier experiments, that has been supplemented by the module EventCalc that samples events similar to traditional Monte Carlo generators.
@@ I Introduction
In this paper, we address these issues by systematically calculating the reach of accelerator-based experiments for long-lived dark photons in the GeV mass range. We incorporate recent advancements in dark photon phenomenology: Ref. Foroughi-Abari:2024xlj , which reanalyzed proton bremsstrahlung production, and forthcoming work Kyselov:2024ta , which investigates production via mixing within the fragmentation chains in deep inelastic collisions. We then implement this updated phenomenology in SensCalc Ovchynnikov:2023cry , a tool designed to calculate event rates for LLPs across various experimental setups in a unified framework. It accounts for the acceptance of LLP decay products in the detector and applies various selection criteria based on event topology, geometric parameters, and energy. Using this, we “recalibrate” the parameter space for dark photons and evaluate the effect of bremsstrahlung uncertainties on the experimental reach.
The paper is organized as follows. In Sec. II, we provide a comprehensive discussion of dark photon phenomenology, including production and decay modes, and parameterize the uncertainties in proton bremsstrahlung. Sec. III describes the methods used to calculate the experimental reach, including the implementation in SensCalc. In Sec. IV, we present recalculated constraints and sensitivities for various experiments, showing how the revised production channels, particularly uncertainties in bremsstrahlung, affect these results. Finally, Sec. V contains our conclusions.
@@ II.1 Production modes
Regarding production via mixing, we implement this mechanism directly in PYTHIA8 Bierlich:2022pfr by replacing the intermediate mesons $V^{0}$ with dark photons at the final stage of the fragmentation chain. This approach allows for a systematic calculation of the total yield and kinematics of long-lived particles (LLPs) as a function of their mass. Such a feature was absent in previous studies Berlin:2018jbm ; Jerhot:2022chi , which approximated the LLP flux by multiplying the meson production cross-section with the squared mixing angle and either assumed identical kinematics for the LLPs and their parent mesons or used unclear relationships. More details on this implementation will be provided in a forthcoming paper Kyselov:2024ta . The primary uncertainty in this channel stems from variations in the PYTHIA setup, which needs to be tuned to match accelerator data. For the far-forward LHC experiments, we adopted the tune from Fieg:2023kld , finding that the flux varies by about 30%, largely independent of the LLP mass. However, detailed tuning studies for other facilities are currently lacking and are deferred to future work. Also, we do not include the contribution from higher excitations of vector mesons, which may sizeably enhance the flux from mixing at masses $m_{V}\gtrsim 1\text{ GeV}$ (to be done in the future).
Let us address proton bremsstrahlung. [^1] A detailed description of this production mechanism is provided in Appendix A. Consider the cross-section of the scattering of two protons $P,p$ where dark photons are emitted from the proton leg. If concentrating solely on the initial state radiation part of the process, its cross-section may be approximated by the cross-section of the inelastic scattering of two protons $p,p^{\prime}$ times the “splitting probability” $\omega_{\text{spl}}$ describing the virtual transition $P\to p^{\prime}+V$. Another key quantity that defines the bremsstrahlung flux is the elastic proton form-factor. Because of the mixing of $V$ with vector mesons $V^{0}$, it gets resonantly enhanced at the vicinity of $V^{0}$ masses.
@@
Lastly, for SPS and other beam dump experiments, it may be important to consider the cascade production of dark photons. We do not consider it in our analysis. We motivate this for two reasons. First, the secondarily produced $\pi^{0},\eta,\dots$ are much softer than the primary ones, and have a broad angular distribution. These two features significantly decrease their contribution to the flux of $V$s, pointing to the decay volume and having large enough energy to be detected (see a discussion in SHiP:2020sos ; SHiP:2020vbd ). Second, the $e^{+}e^{-}$ annihilation and $e^{\pm}$ bremsstrahlung mainly enhance the low-mass production flux $m_{V}\lesssim 100\text{ MeV}$ in the domain of relatively low energies. The corresponding parameter space is mainly excluded by astrophysical constraints and past experiments.
@@ II.1.1 Production channels hierarchy
Figure 2: Top panel: overall probabilities to produce dark photons (total number of produced dark photons normalized by the number of proton collisions) via various channels described in Sec. II.1. The SPS facility is considered as an example. The blue curves indicate various descriptions of the bremsstrahlung channel: the Baseline description commonly used in the literature (the dashed blue line) and the FR description from the paper Foroughi-Abari:2024xlj (the blue line), with the uncertainties for the latter indicated by the light blue band. The width of the band is caused by the variation of the proton virtuality scale and the proton elastic electromagnetic form factor in the time-like region (see also Appendix A). Bottom panel: the contribution of these production channels to the differential number of events in dark photon’s energy at the SHiP experiment. For the bremsstrahlung, the “Central” description is used. The value of the mixing angle close to the upper bound of the sensitivity is considered (i.e., $c\tau\langle\gamma\rangle$ is smaller than the distance from the target to the decay volume), which highlights the importance of the bremsstrahlung channel even in the mass range where the overall production probability is suppressed. See Sec. III on the method we used to obtain the distribution.
@@ II.2 Decay modes
The branching ratios of various decays, as well as the resulting proper lifetime of dark photons, $c\tau_{V}$, are shown in Fig. 3. Leptonic modes dominate in the mass range $m_{V}\lesssim 0.5\text{ GeV}$. The hadronic decays have complicated scaling with mass because of the mixing with $\rho^{0},\omega,\phi$. It leads to the resonant enhancement of the probabilities of the corresponding decays, such as $V\to\pi^{+}\pi^{-}$ (the $\rho^{0}$-like decay), $V\to 2K$ (the $\phi$-like decay), and $V\to 3\pi$ (the $\omega$-like decay). The inclusive and exclusive descriptions match at mass $m_{V}\approx 1.6\text{ GeV}$.
@@ III Approach to calculate the reach of experiments
The key experiments include: the past NuCal Blumlein:1990ay , BEBC BEBCWA66:1986err ; WA66:1985mfx , and CHARM CHARM:1983ayi ; CHARM:1985anb ; the currently running NA62 NA62:2023qyn ; Antel:2023hkf ; the approved SHiP Alekhin:2015byh ; Albanese:2878604 and DarkQuest Batell:2020vqn ; Apyan:2022tsd ; and the proposed FASER2 Kling:2021fwx in its FPF configuration (FASER2-FPF).
In this analysis, we do not consider some of the currently running and future experiments, including FASER, SND@LHC, FACET Cerci:2021nlb , the Downstream algorithm at LHC, and DUNE. Despite being the on-axis experiments, the sensitivity of FASER and SND@LHC is constrained by the parameter space where meson decay dominates – a well-understood regime. This limitation is due to their relatively small geometric acceptance and low integrated luminosity. As such, these experiments do not warrant further revision for this particular study. As for FACET, the Downstream algorithm, and DUNE, their sensitivities may be obtained in a way similar to the sensitivities of the other future experiments (see the section below).
@@ III.1 Approach to calculate the sensitivities
To calculate the reach of various experiments in a unified fashion, we will use the SensCalc package Ovchynnikov:2023cry , [^2] May be found on Zenodo and Github which calculates the number of events at various experiments using the semi-analytic approach. It includes all the experiments of interest and the production and decay modes of dark photons discussed in Sec. II.
@@
For all the experiments, we obtain the sensitivity curves by requiring $N_{\text{events}}\geq 2.3$, which corresponds to 90% CL, assuming that they were or are expected to be background-free. We start with past and ongoing experiments and then consider future experiments. We will show their reach based on the approach of Foroughi-Abari:2024xlj for the bremsstrahlung and Kyselov:2024ta for the production via mixing. For constraints of the experiments that will not be revised in this study, we will use the curves from Antel:2023hkf , together with added constraints from NA62 in the dump mode from NA62:2023qyn and FASER FASER:2023tle . For the NA62 curve, the bremsstrahlung already sizeably contributes to the excluded region. Given that bremsstrahlung already contributes to the NA62 constraints and we cannot reproduce it, we have not included it.
@@ V Conclusions
Using this machinery, we have carefully analyzed how the constraints and sensitivities of various experiments change depending on the variation of the proton bremsstrahlung flux – the main source of uncertainty in the event rate (Sec. II), which may change it by orders of magnitude, see Sec. IV. We have found that this translates into varying the maximal mass probed by the given experiment by as large as a factor of two, depending on the setup, both for past (Sec. IV.1) and future (Sec. IV.2) experiments. To demonstrate the magnitude of the uncertainties, Fig. 6 shows the parameter space of dark photons where the proton bremsstrahlung flux is considered at its lower and upper values within the uncertainty band.
In the end, let us highlight the directions to be explored. First, the revision of the dark photon parameter space performed in this work is far from being complete. Constraints and sensitivities from the other probes, such as prompt searches at the LHC, may rely on the dark photon bremsstrahlung and thus have to be recomputed. Dark photons are not the only well-motivated model of LLPs where the theoretical uncertainties in the production and decay modes may severely affect the status of the parameter space. The phenomenology of Higgs-like scalars, $B-L$ mediators, axion-like particles, and other LLPs has to be carefully analyzed. This paper indicates the unified and systematic approach to how this may be done. We leave these interesting questions for future work.

### v3

@@ Abstract
A systematic and unified analysis of the capability of lifetime frontier experiments to probe the parameter space of hypothetical long-lived particles is crucial for defining the targets of future searches. However, such a comprehensive study has not yet been conducted for dark photons – hypothetical massive particles that kinetically mix with Standard Model photons. Existing studies often rely on outdated assumptions about dark photon phenomenology, leading to inaccurate calculations of the parameter space, including regions already excluded or relevant for future experiments. In this paper, we present a comprehensive calculation of the parameter space for GeV-scale dark photons and light dark matter accessible to lifetime frontier experiments. Our analysis highlights how theoretical uncertainties in dark photon phenomenology can significantly impact experimental sensitivity across coupling and mass parameters. To facilitate further investigation, we provide these results in a publicly accessible form, incorporating updated dark photon phenomenology into the event generator SensCalc.
@@ I Introduction
In this paper, we address these issues by systematically calculating the reach of accelerator-based experiments for long-lived dark photons in the GeV mass range. We incorporate recent advancements in dark photon phenomenology: Ref. Foroughi-Abari:2024xlj , which reanalyzed proton bremsstrahlung production, and forthcoming work Kyselov:2025ta , which investigates production via mixing within the fragmentation chains in deep inelastic collisions. We then implement this updated phenomenology in SensCalc Ovchynnikov:2023cry , a tool designed to calculate event rates for LLPs across various experimental setups in a unified framework. It accounts for the acceptance of LLP decay products in the detector and applies various selection criteria based on event topology, geometric parameters, and energy. Using this, we “recalibrate” the parameter space for dark photons and evaluate the effect of bremsstrahlung uncertainties on the experimental reach.
The paper is organized as follows. In Sec. II, we provide a comprehensive discussion of dark photon phenomenology, including production and decay modes, and parameterize the uncertainties in proton bremsstrahlung. Sec. III describes the methods used to calculate the experimental reach, including the implementation in SensCalc. In Sec. IV, we present recalculated constraints and sensitivities for various experiments, showing how the revised production channels, particularly uncertainties in bremsstrahlung, affect these results. There, we also consider the case of the elastic dark matter coupled to dark photons – another widely considered model of new physics. Finally, Sec. V contains our conclusions.
@@ II.1 Production modes
Regarding the production via mixing, we implement this mechanism directly in PYTHIA8 Bierlich:2022pfr by replacing the intermediate mesons $V^{0}$ with dark photons at the final stage of the fragmentation chain. This approach allows for a systematic calculation of the total yield and kinematics of long-lived particles (LLPs) as a function of their mass. Such a feature was absent in previous studies Berlin:2018jbm ; Jerhot:2022chi , which approximated the LLP flux by multiplying the meson production cross-section with the squared mixing angle and either assumed identical kinematics for the LLPs and their parent mesons or used unclear relationships.
More details on this implementation will be provided in a forthcoming paper Kyselov:2025ta . The primary uncertainty in this channel stems from variations in the PYTHIA setup, which needs to be tuned to match accelerator data. For the far-forward LHC experiments, we adopted the tune from Fieg:2023kld , finding that the flux varies by about 30%, largely independent of the LLP mass. Detailed tuning studies for other facilities are currently lacking and are deferred to future work. However, following the LHC case, we do not expect the variation of the uncertainty within the range beyond $\mathcal{O}(1)$. Also, we do not include the contribution from higher excitations of vector mesons, which may sizeably enhance the flux from mixing at masses $m_{V}\gtrsim 1\text{ GeV}$. The reason is that the production of such mesons in the fragmentation chain is not incorporated in PYTHIA8. Nevertheless, the resulting flux is a conservative estimate.
Let us switch to the proton bremsstrahlung. [^1] A detailed description of this production mechanism is provided in Appendix A. Consider the cross-section of the scattering of two protons $P,p$ where dark photons are emitted from the proton leg. If concentrating solely on the initial state radiation part of the process, its cross-section may be approximated by the cross-section of the inelastic scattering of two protons $p,p^{\prime}$ times the “splitting probability” $\omega_{\text{spl}}$ describing the virtual transition $P\to p^{\prime}+V$. Another key quantity that defines the bremsstrahlung flux is the elastic proton form-factor. Because of the mixing of $V$ with vector mesons $V^{0}$, it gets resonantly enhanced at the vicinity of $V^{0}$ masses.
@@
[…]
Lastly, for SPS and other beam dump experiments, it may be important to consider the cascade production of dark photons. We do not incorporate this in our analysis. We motivate this for two reasons. First, the secondarily produced $\pi^{0},\eta,\dots$ are much softer than the primary ones, and have a broad angular distribution. These two features significantly decrease their contribution to the flux of $V$s, pointing to the decay volume and having large enough energy to be detected (see a discussion in SHiP:2020sos ; SHiP:2020vbd ). Second, the $e^{+}e^{-}$ annihilation and $e^{\pm}$ bremsstrahlung mainly enhance the low-mass production flux $m_{V}\lesssim 100\text{ MeV}$ in the domain of relatively low energies. The corresponding parameter space is mainly excluded by astrophysical constraints and past experiments.
@@ II.1.1 Production channels hierarchy
Figure 2: Top panel: overall probabilities to produce dark photons (total number of produced dark photons normalized by the number of proton collisions) via various channels described in Sec. II.1. The SPS facility is considered as an example. The blue curves indicate various descriptions of the bremsstrahlung channel: the Baseline description commonly used in the literature (the dashed blue line) and the FR description from the paper Foroughi-Abari:2024xlj (the blue line), with the uncertainties for the latter indicated by the light blue band. The width of the band is caused by the variation of the proton virtuality scale and the proton elastic electromagnetic form factor in the time-like region (see also Appendix A). The red domain originates from the uncertainty in the description of the Drell-Yan process. Bottom panel: the contribution of these production channels to the differential number of events in dark photon’s energy at the SHiP experiment. For the bremsstrahlung, the “Central” description is used. The value of the mixing angle close to the upper bound of the sensitivity is considered (i.e., $c\tau\langle\gamma\rangle$ is smaller than the distance from the target to the decay volume), which highlights the importance of the bremsstrahlung channel even in the mass range where the overall production probability is suppressed. See Sec. III on the method we used to obtain the distribution.
@@ II.2 Decay modes
The branching ratios of various decays, as well as the resulting proper lifetime of dark photons, $c\tau_{V}$, are shown in Fig. 3. Leptonic modes dominate in the mass range $m_{V}\lesssim 0.5\text{ GeV}$. The hadronic decays have complicated scaling with mass because of the mixing with $\rho^{0},\omega,\phi$. It leads to the resonant enhancement of the probabilities of the corresponding decays, such as $V\to\pi^{+}\pi^{-}$ (the $\rho^{0}$-like decay), $V\to 2K$ (the $\phi$-like decay), and $V\to 3\pi$ (the $\omega$-like decay).
The inclusive and exclusive descriptions match at mass $m_{V}\approx 1.6\text{ GeV}$. Overall, the theoretical uncertainty coming from this description of dark photon decays is around $20\%$ in the given mass range Ilten:2018crw .
@@ III Approach to calculate the reach of experiments
The experiments we consider are: the past NuCal Blumlein:1990ay , BEBC BEBCWA66:1986err ; WA66:1985mfx , and CHARM CHARM:1983ayi ; CHARM:1985anb ; the currently running NA62 NA62:2023qyn ; Antel:2023hkf ; the approved SHiP Alekhin:2015byh ; Albanese:2878604 and DarkQuest Batell:2020vqn ; Apyan:2022tsd ; and the proposed FASER2 Kling:2021fwx in its FPF configuration (FASER2-FPF).
In this analysis, we do not consider some of the currently running and future experiments, including FASER, SND@LHC, FACET Cerci:2021nlb , the Downstream algorithm at LHCb Gorkavenko:2023nbk , and DUNE. Despite being the on-axis experiments, the sensitivity of FASER and SND@LHC is constrained by the parameter space where meson decay dominates – a well-understood regime. This limitation is due to their relatively small geometric acceptance and low integrated luminosity. As such, these experiments do not warrant further revision for this particular study. As for FACET, the Downstream algorithm, and DUNE, their sensitivities may be obtained in a way similar to the sensitivities of the other future experiments (see the section below), and we do not include them in order not to overfill the final plots.
Also, we do not recalculate the constraints from astrophysical probes from supernova, and prompt searches from LHCb Ilten:2016tkc . In the supernova environment, protons are non-relativistic, and the main production of dark photons is the quasi-elastic bremsstrahlung. Therefore, the flux of dark photons may be accurately calculated without introducing the virtuality form factor, see Ref. Rrapaj:2015wgs , and the constraint is unaffected by uncertainties. Regarding LHCb, the re-analysis would require a detailed background calculation, which is only possible using the full LHCb simulation framework and goes beyond the scope of this work.
When incorporating the uncertainties in the dark photon phenomenology, we only include those above the $\mathcal{O}(1)$ range: from the proton bremsstrahlung and the Drell-Yan process. It is meaningful because the much smaller uncertainties from the other production modes are comparable with other possible uncertainty sources that we do not include in our work (which are experiment-dependent and include the variation of the reconstruction efficiencies, effects of the finite target, etc.). Moreover, unlike the uncertainties in the bremsstrahlung and Drell-Yan process, they are invisible in the parameter space in the logarithmic scale.
@@ III.1 Approach to calculate the sensitivities
To calculate the reach of various experiments in a unified fashion, we will use the SensCalc package Ovchynnikov:2023cry , [^2] May be found on Zenodo SensCalc-Zenodo and Github which calculates the number of events at various experiments using the semi-analytic approach. It includes all the experiments of interest and the production and decay modes of dark photons discussed in Sec. II.333 Ref. Foroughi-Abari:2024xlj also presents the impact of the proton bremsstrahlung uncertainty on the reach of some experiments. However, only the future experiments SHiP and FASER2 are considered. In addition, Ref. Foroughi-Abari:2024xlj uses FORESEE package, which does not calculate the effects of the decay products acceptance – an essential feature for beam dump experiments that significantly affect the number of events Bondarenko:2023fex .
@@
For all the experiments, we obtain the sensitivity curves by requiring $N_{\text{events}}\geq 2.3$, which corresponds to 90% CL, assuming that they were or are expected to be background-free. We start with past and ongoing experiments and then consider future experiments. We will show their reach based on the approach of Foroughi-Abari:2024xlj for the bremsstrahlung and Kyselov:2025ta for the production via mixing. For constraints of the experiments that will not be revised in this study, we will use the curves from Antel:2023hkf , together with added constraints from NA62 in the dump mode from NA62:2023qyn and FASER FASER:2023tle . For the NA62 curve, the bremsstrahlung already sizeably contributes to the excluded region. Given that bremsstrahlung already contributes to the NA62 constraints and we cannot reproduce it, we have not included it.
@@ IV.2 Future experiments
IV.3 Light dark matter
Apart from being a standalone new physics particle, dark photons may serve as a mediator between the Standard Model and hypothetical light dark matter (LDM) $\chi$ Beacham:2019nyx . Since the main production mechanism of $\chi$s is the decay of dark photons into $\chi\bar{\chi}$, the uncertainties in the dark photon phenomenology straightforwardly translate to the parameter space of LDM. There is a notable difference in simplifying the re-analysis: the main constraints on the $\chi$s parameter space come from the electron beam dump experiment NA64 NA64:2023wbi and the electron-positron collider BaBar BaBar:2017tiz . The production modes of dark photons there are cleaner and unaffected by the uncertainties discussed in this study.
Figure 6: Sensitivity of SHiP to the model of the elastic light dark matter $\chi$ coupled via dark photons Beacham:2019nyx ; Antel:2023hkf . Uncertainties in the SHiP sensitivity originate from the dark photon production uncertainty. The constraints are taken from NA64:2023wbi , the projection of the future projection of NA64 from Gninenko:2019qiv , and the thermal relic target line for the pseudo-Dirac $\chi$
The main future experiments probing the parameter space of this model are NA64 Gninenko:2019qiv (with the increased luminosity) and SHiP Antel:2023hkf . SHiP will search for the events with $\chi$s via the elastic scatterings of $\chi$ off electrons in the Scattering and Neutrinos Detector (called SND@SHiP) SHiP:2020noy . To re-evaluate its sensitivity, we use the results of the recent study Ferrillo:2023hhg that performed background analysis for the new SND@SHiP setup. Instead of rescaling the sensitivity curve from SHiP:2020noy (as it was done in Ferrillo:2023hhg ), we recompute the number of events from scratch.
The parameter space of LDM, including the sensitivities of SHiP and future searches at NA64, are shown in Fig. 6. The dark photon uncertainty shows up depending on the ratio between $m_{\chi}$ and $m_{V}$, which we assume to be $1/3$, following the standard convention. From the figure, we see that the uncertainty affects a huge fraction of the SHiP’s target parameter space.
@@ V Conclusions
Using this machinery, we have carefully analyzed how the constraints and sensitivities of various experiments change depending on the variation of the proton bremsstrahlung flux – the main source of uncertainty in the event rate (Sec. II), which may change it by orders of magnitude, see Sec. IV. We have found that this translates into varying the maximal mass probed by the given experiment by as large as a factor of two, depending on the setup, both for past (Sec. IV.1) and future (Sec. IV.2) experiments.
To demonstrate the magnitude of the uncertainties, Fig. 7 shows the parameter space of dark photons where the proton bremsstrahlung flux is considered at its lower and upper values within the uncertainty band.
Our findings are significant for several reasons. First, recognizing that the constraints from past experiments are heavily influenced by theoretical uncertainties reshapes the target parameter space for future searches. Second, as the sensitivity of upcoming experiments, such as FASER2, varies considerably, their physics case and overall appeal require careful reassessment. Third, we have identified a critical limitation in the state-of-the-art studies Beacham:2019nyx ; Antel:2023hkf : the lack of unified recommendations for the experimental community regarding the phenomenology of long-lived particles. This inconsistency leads to incoherent presentations of constraints and sensitivities across various experiments in summary figures. Addressing this issue promptly is essential to fully assess the potential of future experiments to explore the parameter space.
In the end, let us highlight the directions to be explored. First, the revision of the dark photon parameter space performed in this work is far from being complete. Constraints and sensitivities from the other probes, such as prompt searches at the LHC, may rely on the dark photon bremsstrahlung and thus have to be recomputed.
Dark photons are not the only well-motivated model of LLPs where the theoretical uncertainties in the production and decay modes may severely affect the status of the parameter space. The impact of the theoretical uncertainty in the phenomenology on the probed parameter space for Higgs-like scalars, $B-L$ mediators, axion-like particles, and other LLPs has to be carefully analyzed. However, these models also have uncertainties coming from the production via the proton bremsstrahlung and mixing with neutral mesons Monin:2018lee ; Blackstone:2024ouf ; Smith:2024jve . Therefore, these studies are a subject of dedicated research. This paper indicates the unified and systematic approach to how this may be done. We leave these interesting questions for future work.